OpenStudy (anonymous):

if z=2cos 60degree, find Z^4 in retangular form.

OpenStudy (anonymous):

@Hero , do you ind explaining with a formula?

OpenStudy (kohai):

So an easy way to do this is to just calculate 2cos60, and then multiply itself by that four times. So like, \[1^4 = 1 \times 1 \times 1 \times 1\]

OpenStudy (blank ):

That's what I thought too. ^

OpenStudy (anonymous):

Which is 1, times z^4 would be 1.

OpenStudy (kohai):

Well, 1 was an example to show you what you would be doing. Z would be the result of 2cos60

OpenStudy (anonymous):

I can show options.

OpenStudy (kohai):

Okay here, first off, what is 2cos60?

OpenStudy (anonymous):

It's 1.

OpenStudy (kohai):

Oh, ew. Okay, so that's what it's asking

OpenStudy (kohai):

Are you sure that's supposed to say cis not cos?

OpenStudy (anonymous):

It's cos.

OpenStudy (anonymous):

Which is still equal to 1.

OpenStudy (kohai):

Right, yeah, I know. I honestly don't know how to do this with i numbers

OpenStudy (anonymous):

Thank you for trying.

OpenStudy (kohai):

Sorry I couldn't be of more help

OpenStudy (anonymous):

I'll see if someone else can help.